plese help me out with this. Im totally confused on what to do. Heres the problem- An express train travels 80km per hour from Ironton to Wildwood. A local train, traveling at 48km per hour, takes 2 hours longer for the same trip. How far apart are Ironton and Wildwood?
Found 2 solutions by tutorcecilia, ankor@dixie-net.com:
Answer by tutorcecilia(2152)

Use the distance formula:
Distance = (Rate)(Time)
.
Express: Distance = (80kph)(time)
Local: Distance = (48kpm)(time + 2 additional hours)
Find the distance apart.
.
Since the both trains will travel the same distance between the towns:
Distance (Express)=Distance (Local)
80t=(48)(t+2) [Simplify]
80t=48t+96
80t-48t=96
32t=96
t=96/32
t=3 hours
.
Plug 3 hours back into either of the original formulas and solve for the distance.
Express: Distance=(80)(time)
Distance = (80)(3)
Distance = 240 miles.
.
Try the other distance for the local train:
Local: Distance = (48)(time + 2 additional hours)
Distance = (48)(3+2)
Distance = (48)(5)
Distance = 240 miles

Answer by ankor@dixie-net.com(22740)

An express train travels 80km per hour from Ironton to Wildwood. A local train, traveling at 48 km per hour, takes 2 hours longer for the same trip. How far apart are Ironton and Wildwood?
:
If we find the time that it takes for each train to make the trip it is an easy calculation to find the distance;
:
Let t = time for the express train to make the trip
Then (t+2) = time for the local to make the same trip
:
Dist = speed * time:
Express distance = 80t
Local distance = 48(t+2)

We know that the distance the two trains travel is the same, a simple equation:
:
Express dist = Local dist
80t = 48(t+2)
80t = 48t + 96
80t - 48t = 96
32t = 96
t = 96/32
t = 3 hrs for the express
:
Of course the local would take 5 hrs
:
Remember the Express distance eq, 80t: 80 * 3 = 240 miles
:
If we did it correctly, the local distance eq should be the same: 48 * 5 = 240 also.
:
Did this unconfuse you somewhat??